Let f(t)=t^2+4t+2.
Find a value of t such that the average rate of change of f(t) from 0 to t equals 13

Answer by nerdybill(6962)

Let f(t)=t^2+4t+2.
Find a value of t such that the average rate of change of f(t) from 0 to t equals 13
.
"average rate of change" is simply the instantaneous slope when t=13.
.
To find it, you can find the derivative of f(t):
f(t)=t^2+4t+2
f'(t)=2t+4
.
Now, find f'(13):
f'(t)=2t+4
f'(13)=2(13)+4
f'(13)=26+4
f'(13)=30